Your client weighs 75 kg. Mix 250 units of Humulin regular insulin in 500 mL of...

  1. Your client weighs 75 kg. Mix 250 units of Humulin regular insulin in 500 mL of 0.9% NS. 2) Begin continuous Humulin regular insulin infusion at 0.1 units/kg/hour IV. 3) Calculate _____units = _____mL/hr            
Answer #1

ANSWER.

* Order = 0.1 unit/kg/hour.

   Formula = Desired dose / Dose in hand * Quantity.

       Weight of the patient = 75 kg.

     So

       Desired dose(D) = 75 * 0.1

                             = 7.5 units.

       Dose in hand (H) = 250 unit.

            Quantity (Q) = 500 ml.

As per formula

        Amount (units) of insulin needed = D/H * Q

                                                             = 7.5/250 * 500

                                                             = 0.03 * 500

                                                             = 15 ml

      Amount required as ml/hr = 15 ml/hr.

      So

               * 7.5 units = 15 ml/hr.